A stretched wire (fixed at both ends) is connected to a vibrator with fixed frequency f. Under a tension, T, a standing wave pattern consisting of 5 loops is observed on the wire. What should be the new tension T’ in order to obtain one loop on the wire?


1
Expert's answer
2021-06-01T10:50:30-0400
"f=\frac{1}{2\pi}\sqrt{\frac{T}{\mu}}\\T'=5^2T=25T"